18.2 Two-level cache performance A processor with two levels of caches has a CPI of 1 when there is no level-1 cache miss. At level 1, the
media%2F4ea%2F4eadc871-f1cd-4c72-ba5e-e4
Answer #1

a) effective cpi= 0.95*1 +0.05*(0.98*10+0.02*60)=0.95+0.05*11=1.50

b) 1.50=p*1+(1-p)*10

     10-9*p=1.5

     p=8.5/9=0.94444444

     1-p=0.05555555